OpenStudy (klepez):

Simplify. 2 over the square root of 5 (1 point)

OpenStudy (klepez):

here are the choices \[\sqrt{10}\]

OpenStudy (klepez):

\[2\sqrt{5}\]

OpenStudy (klepez):

\[\frac{ \sqrt{10} }{ 5 }\]

OpenStudy (klepez):

\[\frac{ 2\sqrt{5} }{ 5 }\]

OpenStudy (lexipoo1998):

I think it is the last one but i might be wrong...

OpenStudy (klepez):

ok thanks

OpenStudy (lexipoo1998):

No problem

OpenStudy (raden):